OpenStudy (anonymous):

Find the distance between the given points. C(7, 1) and D(4, -6) Distance =

OpenStudy (anonymous):

49+9 58 radical58

OpenStudy (anonymous):

man ur smart

OpenStudy (anonymous):

The straight line distance would be given by use of pythagoras: Distance=sqrt((change in y)^2 + (change in x)^2) which in this case; Distance=sqrt(7^2+3^2)=sqrt(58)